# The quantum Newton’s bucket: Active and passive rotations in quantum theory

### v1

@@ V Conclusions
This work has spanned a large range of topics, ranging from famous scientific texts Newton (1687); van Vleck (1934) to technical problems of current interest in specialized topics such as QCD at finite temperature Braguta et al. (2023); Chernodub (2021); Ayala et al. (2021); Prokhorov et al. (2020); Yamamoto and Hirono (2013); Aristova et al. (2016); Wang et al. (2019). As we have seen, the problem of whether passive and active rotations physically coincide in a given quantum system is, as expected, intimately related to the symmetries of the system and of the role symmetry has in quantum mechanics (which can be significantly different from the classical counterpart).
@@
While insights from quantum mechanics can be generalized to quantum field theory by symmetry arguments, ascertaining the quantitative discrepancy, necessary for phenomenology (is the difference between active and passive rotations a small correction or a large one in a given system?), is a much more challenging exercise for which at the moment we can not give even a qualitative recipe (beyond decompositions such as Ji (1997) and operator product expansions of the type used in Kim et al. (2023).
@@
While we concentrated on rotations, the arguments given here can be easily generalized to non-inertial frames. A lot of the calculations done for this case have involved accellerations “put in by hand”. This has led to quite a lot of debates in the literature about whether the Unruh effect “is real”. Understanding general criteria for the differences in passive and active accellerations (the latter with an explicit interaction with a classical field included Brout et al. (1991); Labun and Rafelski (2012); Hegelich et al. (2022)) could be central to this question. In this respect, a clear example of this is provided by the recent extension of the COW experiment Colella et al. (1975) to freely falling wavepackets Dobkowski et al. (2025): Looking at FIg. 1 of Dobkowski et al. (2025) one can think of an ”active” transformation as a detector comoving with the red wavepacket measuring the phase shift of the blue wavepacket, and a ”passive” transformation as the opposite’ (the detector is assumed to be performing measurements in intervals longer than the width of the wavepacket). The equality of the two phase shifts is a good illustration that in gravitational dynamics active and passive transformations are indeed the same.
@@
Beyond heavy ion physics, the sort of questions examined in this paper are of relevance for the rapidly growing field of quantum reference frames Giacomini et al. (2019); Toroš et al. (2017); Padmanabhan and Padmanabhan (2011) and their relationship to the quantization of gravity. Ultimately, the arguments around Newton’s bucket were satisfactorily settled by Einstein, who worked out that it is the local gravitational field (produced by the bucket, the water, the earth etc.) that determines the family of local “inertial” coordinate systems which determine which set of transformations are equivalent between active and passive. Of course this is still the last theory that is resistent to quantization. Perhaps, the theoretical Giacomini et al. (2019); Toroš et al. (2017) and experimental Colella et al. (1975); Rauch and Werner (2015) study of quantum systems in non-inertial reference frames could give clues Torrieri (2023, 2024b) to the right procedure to accomplish this task.

### v2

@@ 5 Conclusions
This work has spanned a large range of topics, ranging from famous scientific texts [Newton(1687), van Vleck(1934)] to technical problems of current interest in specialized topics such as QCD at finite temperature and vorticity [Braguta et al.(2023)Braguta, Kudrov, Roenko, Sychev, and Chernodub, Chernodub(2021), Ayala et al.(2021)Ayala, Hernández, Raya, and Zamora, Prokhorov et al.(2020)Prokhorov, Teryaev, and Zakharov, Yamamoto and Hirono(2013), Aristova et al.(2016)Aristova, Frenklakh, Gorsky, and Kharzeev, Wang et al.(2019)Wang, Wei, Li, and Huang] (and more generally accelleration [Ohsaku(2004), Benic and Fukushima(2015), Salluce et al.(2024)Salluce, Pasini, Flachi, Pittelli, and Ansoldi, Casado-Turrión and Dobado(2019), Chernodub(2025), Unruh and Weiss(1984)], where a controversy of the exact role of the ”Unruh effect” in phase transitions exists in the literature). As we have seen, the problem of whether passive and active rotations physically coincide in a given quantum system is, as expected, intimately related to the symmetries of the system and of the role symmetry has in quantum mechanics (which can be significantly different from the classical counterpart).
@@
While insights from quantum mechanics can be generalized to quantum field theory by symmetry arguments, ascertaining the quantitative discrepancy, necessary for phenomenology (is the difference between active and passive rotations a small correction or a large one in a given system?), is a much more challenging exercise for which at the moment we can not give even a qualitative recipe (beyond decompositions such as [Ji(1997)] and operator product expansions of the type used in [Kim et al.(2023)Kim, Cho, and Lee, Kim et al.(2025)Kim, Gubler, and Sasaki].
While we concentrated on rotations, the arguments given here can be easily generalized to non-inertial frames. A lot of the calculations done for this case have involved accellerations “put in by hand”. This has led to quite a lot of debates in the literature about whether the Unruh effect “is real”. Note that as shown in [Dunajski(2023)] the Unruh effect has a non-relativistic remnant, making it amenable to the sort of quantum mechanical analysis performed in this work. Mixed particle states, allowed in quantum field theory, yield additional issues, particularly if the mixing is between mass and charge states (as in neutrinoes) rather than between two types of charges (as in quarks): As shown in [Ahluwalia et al.(2016a)Ahluwalia, Labun, and Torrieri, Ahluwalia et al.(2016b)Ahluwalia, Labun, and Torrieri] a vacuum of mass states will yield an automatic violation of passive vs active rotations due to the thermal vacuum being defined in terms of mass states. The alternative, assuming the vacuum to be charged states, will yield a small violation of Lorentz symmetry and a non-thermal Unruh spectrum [Blasone et al.(2023)Blasone, Giacosa, Smaldone, and Torrieri, Blasone et al.(2022)Blasone, Lambiase, Luciano, and Petruzziello].
@@
Understanding general criteria for the differences in passive and active accellerations (the latter with an explicit interaction with a classical field included [Brout et al.(1991)Brout, Parentani, and Spindel, Labun and Rafelski(2012), Hegelich et al.(2022)Hegelich, Labun, Labun, Torrieri, and Truran]) could be central to this question. In this respect, a clear example of this is provided by the recent extension of the COW experiment [Colella et al.(1975)Colella, Overhauser, and Werner] to freely falling wavepackets [Dobkowski et al.(2025)Dobkowski et al.]: Looking at FIg. 1 of [Dobkowski et al.(2025)Dobkowski et al.] one can think of an ”active” transformation as a detector comoving with the red wavepacket measuring the phase shift of the blue wavepacket, and a ”passive” transformation as the opposite’ (the detector is assumed to be performing measurements in intervals longer than the width of the wavepacket). The equality of the two phase shifts is a good illustration that in gravitational dynamics active and passive transformations are indeed the same.
@@
Hence, beyond heavy ion physics, the sort of questions examined in this paper are of relevance for the rapidly growing field of quantum reference frames [Giacomini et al.(2019)Giacomini, Castro-Ruiz, and Brukner, Toroš et al.(2017)Toroš, Großardt, and Bassi, Padmanabhan and Padmanabhan(2011)] and their relationship to the quantization of gravity. Ultimately, the arguments around Newton’s bucket were satisfactorily settled by Einstein, who worked out that it is the local gravitational field (produced by the bucket, the water, the earth etc.) that determines the family of local “inertial” coordinate systems which determine which set of transformations are equivalent between active and passive. Of course this is still the last theory that is resistent to quantization. Perhaps, the theoretical [Giacomini et al.(2019)Giacomini, Castro-Ruiz, and Brukner, Toroš et al.(2017)Toroš, Großardt, and Bassi] and experimental [Colella et al.(1975)Colella, Overhauser, and Werner, Rauch and Werner(2015)] study of quantum systems in non-inertial reference frames could give clues [Torrieri(2023), Torrieri(2024b)] to the right procedure to accomplish this task.